The mean of binomial distribution is and the variance is 9/4.Evaluate PMF


1
Expert's answer
2021-06-21T16:35:53-0400

If "X\sim Bin(n, p)," then "E[X]=np, V[X]=np(1-p), \sigma_X=\sqrt{np(1-p)}."

Given "E[X]=3, V[X]=\dfrac{9}{4}." Then


"np=3"

"np(1-p)=\dfrac{9}{4}"

"1-p=\dfrac{3}{4}"

"p=\dfrac{1}{4}"

"n=12"

"f(x; 12,\dfrac{1}{4})= \begin{cases}
 \dbinom{12}{x}(\dfrac{1}{4})^x(1-\dfrac{1}{4})^{12-x},x=0, 1, ..., 12 \\
 otherwise
\end{cases}"